Determine the density of oxygen gas (O2) at a pressure of 172,000Pa and temperature =-18 Celsius. State your answer to the nearest 0.001 kg/m3, follow the standard rounding rules in stating answer. Take the molar mass of molecular oxygen to be 32 g/mol, R=8.31 JK^-1mol^-1 and take the absolute zero to be -273 Celsius.


1
Expert's answer
2021-10-03T13:24:54-0400

Write the universal gas law:


"PV=\frac m{\Mu} RT,\\\space\\
P=\frac{\rho }{\Mu}RT,\\\space\\
\rho=\frac{P\Mu}{RT}=\frac{172000·0.032}{8.31·(273-18)}=2.596\text{ kg/m}^3."